# Structure-Fair Quantum Circuit Complexity: An Auditable Information-Theoretic Lower Bound

### v2

@@ 3.1 Main Theorem
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\geq\frac{1}{\log\Gamma_{R}}D(\rho||\sigma_{R})-\frac{c_{1}}{\log\Gamma_{R}}\left[\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\})+\log(1/\varepsilon)\right]+\frac{1}{\log\Gamma_{R}}\Delta_{\mathrm{spec}}(\rho)$ (3.1)
@@
$\Delta_{\mathrm{spec}}(\rho)=S(\rho)-H_{\mathrm{min}}(\rho)$ (3.2)
where $S(\rho)$ is the von Neumann entropy, and $H_{\mathrm{min}}(\rho)=-\log\lVert\rho\rVert_{\infty}$ is the min-entropy, determined by the largest eigenvalue of $\rho$.Eq. (3.1) constitutes the theoretical core of this paper. The right-hand side of the inequality consists of three terms with clear physical roles:
@@
    Linear Leading Term: $\frac{1}{\log\Gamma_{R}}D(\rho||\sigma_{R})$, which is $C_{R}(\rho)$. It profoundly reveals that the minimum number of computational steps required to prepare a quantum state is, fundamentally, determined by the structurally fair complexity of that state. The more non-trivial information a state contains within its structured vacuum $\sigma_{R}$ (the reference state), the higher the computational cost, in a linear relationship, that any algorithm preparing it must incur.
@@
(Note: It is worth noting that Eq. (3.1) is a bootstrap inequality. Its non-recursive, explicit solution (a closed-form lower bound), which is convenient for direct use in data analysis, is derived in detail along with an implementation scheme in Appendix C.)
3.2 Proof Blueprint
@@
    Algorithmic Information Theory Bound ($K_{\varepsilon}(\rho|R)\to D_{\max}^{\varepsilon}$): A fundamental result in quantum algorithmic information theory is that the smoothed Kolmogorov complexity of a quantum state, $K_{\varepsilon}(\rho|R)$, is lower-bounded by its smoothed max-relative entropy $D_{\max}^{\varepsilon}(\rho||\sigma_{R})$ with respect to the reference state $\sigma_{R}$ [23, 24, 25, 26, 27]. This step connects an uncomputable quantity, $K_{\varepsilon}$, with a rigorously defined information-theoretic quantity, $D_{\max}^{\varepsilon}$. (Note: In particular, a key technical point is that the universal constant in this bound is independent of the system’s dimension, thereby ensuring the universality of our final lower bound. Its rigorous proof can be found in Appendix B.2.)
@@
3.3 Smoothed Theorem and Operational Connection
@@
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\geq\frac{1}{\log\Gamma_{R}}D_{H}^{\eta}(\rho||\sigma_{R})-\frac{c^{\prime}_{1}}{\log\Gamma_{R}}\left[\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\})+\log(1/\varepsilon)\right]-\frac{c^{\prime}_{2}}{\log\Gamma_{R}}\log(1/\eta)$ (3.3)
@@
The core significance of Theorem 3.2 lies in the fact that it connects the lower bound of complexity to a physical task with the most direct operational definition: distinguishing the quantum state $\rho$ from its structured vacuum (the reference state) $\sigma_{R}$ with an optimal strategy. The fundamental difficulty of this task is quantified by the hypothesis testing relative entropy, $D_{H}^{\eta}(\rho||\sigma_{R})$. This quantity precisely characterizes the minimum achievable exponential decay rate of the type-II error probability, under the condition that the type-I error (mistaking $\sigma_{R}$ for $\rho$) probability does not exceed $\eta$. To rigorously establish this connection, we provide a complete proof in Appendix A.6. This proof systematically employs core tools from one-shot information theory, such as the smoothed max-relative entropy, and ultimately unifies the physical cost of a computational process, $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$, with the statistical limit of an information-theoretic task, $D_{H}^{\eta}(\rho||\sigma_{R})$, through a rigorous chain of inequalities.[21] This ultimately proclaims that RCC is not only a self-consistent theoretical construct but also a quantity that can, in principle, be probed through physical measurements. This theorem not only enhances the robustness of our theoretical framework, but more importantly, it paves a solid theoretical path for the operational schemes proposed in Chapter 4 for estimating the complexity lower bound based on statistical experiments. The standard information-theoretic relations connecting the entropy measures used in this theorem and in Theorem 3.1 are provided in Appendix C.
4 Operational Principles and Information Extraction for RCC
@@
The vitality of any profound physical quantity lies not only in its theoretical self-consistency and elegance, but more so in its connection to observables. The preceding sections have established the theoretical framework of Reference-Contingent Complexity and have proven its central role as a lower bound on the optimal gate count. However, if RCC itself could not be probed by any physical means, this lower bound would lose its practical significance. This chapter aims to bridge this gap between theory and reality by systematically elaborating on several fundamental principles for extracting RCC information from a physical system, thereby transforming an abstract theoretical concept into a set of concrete physical measurement protocols with statistical confidence. This chapter will elaborate on these fundamental principles separately, while a practical protocol that integrates them into a unified operational workflow is provided in Appendix B.3.2. On the other hand, this chapter will primarily discuss protocols in ideal finite-dimensional systems, whereas a rigorous framework for robust handling and conservative lower-bound schemes for the unavoidable energy or mode cutoffs when dealing with infinite-dimensional systems (such as quantum field theories) is detailed in Appendix F.
4.1 Principle A: Quantum Hypothesis Testing
@@
From the fundamental perspective of information theory, the distinguishability between two quantum states is characterized by the theory of hypothesis testing. Theorem 3.2 has already revealed that the smoothed lower bound of RCC is directly related to the hypothesis testing relative entropy $D_{H}^{\eta}(\rho||\sigma_{R})$. This provides us with the first, and theoretically most fundamental, principle of information extraction. Its core task can be formulated as a binary quantum hypothesis testing problem: given a single-copy system, determine whether its state satisfies the null hypothesis $H_{0}:\sigma_{R}$ or the alternative hypothesis $H_{1}:\rho$. Any given measurement (described by a POVM $\{T,I-T\}$) will yield a lower bound on $D_{H}^{\eta}(\rho||\sigma_{R})$, which alleviates the stringent requirement of finding the optimal measurement, making this principle feasible in practice. At the implementation level, let us assume $N$ independent and repeated measurements are performed, from which the frequency of the type-II error (mistaking $\rho$ for $\sigma_{R}$) is observed, and its probability $\beta$ is thereby estimated. To ensure statistical rigor, we need to construct a confidence interval for $\beta$. Using standard statistical methods, such as the Hoeffding bound or the more precise Clopper-Pearson interval [28, 29], one can calculate the upper confidence endpoint of $\beta$, denoted $\beta_{U}(\delta)$, at a confidence level of $1-\delta$. (Note: In the chosen binary test $\{T,I-T\}$, it is specified that observing $T$ leads to the decision $H_{1}(\rho)$, while observing ITI-T leads to the decision $H_{0}(\sigma_{R})$. The type-I error is constrained as $\alpha=\mathrm{Tr}[T\cdot\sigma_{R}]\leq\eta$. With a finite sample size, the upper confidence endpoint of $\alpha$, $\alpha_{U}(\delta)$, can be obtained using the Clopper–Pearson (or Wilson) interval, and it should be ensured that $\alpha_{U}(\delta)\leq\eta$; under this constraint, the lower bound of $D_{H}^{\eta}(\rho||\sigma_{R})$ is then assessed based on the upper confidence endpoint of $\beta$, $\beta_{U}(\delta)$.) We then have a statistical lower bound for the leading term of RCC:
@@
$D_{H}^{\eta}(\rho||\sigma_{R})\geq-\log\beta_{U}(\delta)$ (4.1)
This inequality reveals a profound physical picture: the intrinsic complexity of a quantum state $\rho$ is lower-bounded by our ability to confidently distinguish it from its “structured vacuum” $\sigma_{R}$ in an experiment. A highly complex state must have extremely high distinguishability, corresponding to an extremely low type-II error rate, which in turn implies a large lower bound on its complexity. The complete experimental protocol, including the rigorous control of statistical errors, is detailed in Appendix D.1. Of particular importance is that the advantage of the present theoretical framework lies in the fact that this connection between “distinguishability” and “complexity” does not merely remain at the level of physical intuition. It is elevated by Theorem 3.2 to a mathematically provable relationship. It is this theorem, along with its complete derivation based on one-shot information theory in Appendix A.6, that builds the bridge from experimental data (the type-II error rate $\beta$) to the theoretical lower bound ($C_{\mathrm{opt}}$). This rigorous connection ensures that our proposed operational schemes are theoretically completely reliable and self-consistent, successfully transforming an abstract concept of complexity into a measurable physical reality.
4.2 Principle B: Projective Witnesses
@@
$D_{\max}(\rho||\sigma_{R})\geq\log(p\cdot d_{R}/r)$ (4.2)
Proof. If $\rho\leq 2^{\lambda}\cdot\sigma_{R}$, then for a projector $\Pi$ of rank $r$, taking the trace yields $\mathrm{Tr}(\Pi\cdot\rho)\leq 2^{\lambda}\cdot\mathrm{Tr}(\Pi\cdot\sigma_{R})=2^{\lambda}\cdot r/d_{R}$. From $p=\mathrm{Tr}(\Pi\cdot\rho)$, one can solve for $\lambda\geq\log(p\cdot d_{R}/r)$. The lemma is thus proven. In experimental practice, we estimate the value of $p$ through $N$ measurements and calculate its lower confidence endpoint $p_{\mathrm{lower}}(\delta)$ at a confidence level of $1-\delta$. Substituting this statistical lower bound into Eq. (4.2) yields a lower bound for $D_{\max}$ with statistical confidence. The physical meaning of this result is exceptionally clear: if it is found experimentally that the quantum state $\rho$ occupies, with a high probability $p_{\mathrm{lower}}$, a subspace (defined by $\Pi$) of dimension $r$ much smaller than the dimension $d_{R}$ of its containing structured subspace, then this serves as a proof that the information in this state is highly “concentrated” or “ordered.” This high concentration of information implies that the state $\rho$ possesses a very large eigenvalue, and thus $H_{\mathrm{min}}(\rho)$ is small and $D_{\max}$ is large. Therefore, through a simple projective measurement, we have “witnessed” the non-triviality of the state, setting a rigorous lower bound on its complexity. The detailed implementation protocol for this principle, including the statistical estimation of occupation probabilities and confidence level corrections in scenarios with multiple witnesses, is formally articulated in Appendix D.2.
4.3 Principle C: Dephasing Information Bound
The third principle extracts information through a “classicization” process. To obtain a reliable lower bound, we employ the completely dephasing channel $\Delta_{R}$ derived from the reference set $R$, which corresponds to performing a measurement in a basis of rank-1 projectors that simultaneously diagonalizes all projection operators $\{P_{i}\}$ and then discarding the off-diagonal (coherent) information. Physically, this channel corresponds to the strongest decoherence process that leaves the structure $R$ invariant.[21] According to the property of entropy increase in quantum information theory, the von Neumann entropy after the action of this channel is necessarily no less than before, i.e., $S(\Delta_{R}(\rho))\geq S(\rho)$. The dephased state $\Delta_{R}(\rho)$ is a classical probability distribution, and its Shannon entropy is denoted as $H=S(\Delta_{R}(\rho))$. Therefore, we have $H\geq S(\rho)$, which provides an operational and conservative lower bound for RCC:
@@ 4.3 Principle C: Dephasing Information Bound
Experimentally, this corresponds to performing measurements in a classical basis compatible with the reference set $R$, reconstructing the probability distribution, and estimating its Shannon entropy $H$. An upper confidence interval for $H$ can be given through the DKW inequality or bootstrap methods,[30] thereby propagating the statistical uncertainty to the final lower bound of RCC. The detailed implementation protocol for this principle, particularly concerning how to rigorously estimate Shannon entropy from finite samples with bias correction, is presented with its complete statistical framework in Appendix D.3. Furthermore, the three principles proposed in this chapter can be integrated into an operational pipeline with statistical confidence, as detailed in Appendix B.
4.4 An Important Clarification: The Role of Purity

### v3

@@ III.1 Main Theorem
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ $\geq\frac{1}{\log\Gamma_{R}}\,D(\rho\parallel\sigma_{R})$ (6)
$\quad-\frac{c_{1}}{\log\Gamma_{R}}\biggl[\log\!\Bigl(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\}\Bigr)+\log(1/\varepsilon)\biggr]$
$\quad+\frac{1}{\log\Gamma_{R}}\,\Delta_{\mathrm{spec}}(\rho)$
@@
$\Delta_{\mathrm{spec}}(\rho)=S(\rho)-H_{\mathrm{min}}(\rho)$ (7)
where $S(\rho)$ is the von Neumann entropy, and $H_{\mathrm{min}}(\rho)=-\log\lVert\rho\rVert_{\infty}$ is the min-entropy, determined by the largest eigenvalue of $\rho$.Eq. (6) constitutes the theoretical core of this paper. The right-hand side of the inequality consists of three terms with clear physical roles:
@@
    Linear Leading Term: $\frac{1}{\log\Gamma_{R}}D(\rho||\sigma_{R})$, which is $C_{R}(\rho)$. This result indicates that the minimum computational cost to prepare a state is rigorously lower-bounded by its Reference-Contingent Complexity. The more non-trivial information a state contains within its structured vacuum $\sigma_{R}$ (the reference state), the higher the computational cost, in a linear relationship, that any algorithm preparing it must incur.
@@
(Note: It is worth noting that Eq. (6) is a bootstrap inequality. Its non-recursive, explicit solution (a closed-form lower bound), which is convenient for direct use in data analysis, is derived in detail along with an implementation scheme in Appendix C.)
III.2 Proof Blueprint
@@
    Algorithmic Information Theory Bound ($K_{\varepsilon}(\rho|R)\to D_{\max}^{\varepsilon}$): A fundamental result in quantum algorithmic information theory is that the smoothed Kolmogorov complexity of a quantum state, $K_{\varepsilon}(\rho|R)$, is lower-bounded by its smoothed max-relative entropy $D_{\max}^{\varepsilon}(\rho||\sigma_{R})$ with respect to the reference state $\sigma_{R}$ [Gacs2001, berthiaume2001quantum, Datta2009, Tomamichel2012, LiVitanyi2008]. This step connects an uncomputable quantity, $K_{\varepsilon}$, with a rigorously defined information-theoretic quantity, $D_{\max}^{\varepsilon}$. (Note: In particular, a key technical point is that the universal constant in this bound is independent of the system’s dimension, thereby ensuring the universality of our final lower bound. Its rigorous proof can be found in Appendix B.2.)
@@
III.3 Smoothed Theorem and Operational Connection
@@
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ $\geq\frac{1}{\log\Gamma_{R}}\,D_{H}^{\eta}(\rho\parallel\sigma_{R})$ (8)
$\quad-\frac{c^{\prime}_{1}}{\log\Gamma_{R}}\,\biggl[\log\!\Bigl(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\}\Bigr)+\log(1/\varepsilon)\biggr]$
$\quad-\frac{c^{\prime}_{2}}{\log\Gamma_{R}}\,\log(1/\eta)$
@@
The core significance of Theorem 3.2 lies in the fact that it connects the lower bound of complexity to a physical task with the most direct operational definition: distinguishing the quantum state $\rho$ from its structured vacuum (the reference state) $\sigma_{R}$ with an optimal strategy. The fundamental difficulty of this task is quantified by the hypothesis testing relative entropy, $D_{H}^{\eta}(\rho||\sigma_{R})$. This quantity precisely characterizes the minimum achievable exponential decay rate of the type-II error probability, under the condition that the type-I error (mistaking $\sigma_{R}$ for $\rho$) probability does not exceed $\eta$. To rigorously establish this connection, we provide a complete proof in Appendix A.6. This proof systematically employs core tools from one-shot information theory, such as the smoothed max-relative entropy, and ultimately unifies the physical cost of a computational process, $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$, with the statistical limit of an information-theoretic task, $D_{H}^{\eta}(\rho||\sigma_{R})$, through a rigorous chain of inequalities.[HiaiPetz1991] This ultimately proclaims that RCC is not only a self-consistent theoretical construct but also a quantity that can, in principle, be probed through physical measurements. This theorem not only enhances the robustness of our theoretical framework, but more importantly, it paves a solid theoretical path for the operational schemes proposed Sec. IV for estimating the complexity lower bound based on statistical experiments. The standard information-theoretic relations connecting the entropy measures used in this theorem and in Theorem 3.1 are provided in Appendix C.
IV Operational Principles and Information Extraction for RCC
@@
To be physically meaningful, a complexity measure should satisfy theoretical consistency and be estimable through experimental observables. The preceding sections have established the theoretical framework of Reference-Contingent Complexity and have proven its central role as a lower bound on the optimal gate count. However, if RCC itself could not be probed by any physical means, this lower bound would lose its practical significance. This chapter aims to bridge this gap between theory and reality by systematically elaborating on several fundamental principles for extracting RCC information from a physical system, thereby transforming an abstract theoretical concept into a set of concrete physical measurement protocols with statistical confidence. This chapter will elaborate on these fundamental principles separately, while a practical protocol that integrates them into a unified operational workflow is provided in Appendix B.3.2. On the other hand, this chapter will primarily discuss protocols in ideal finite-dimensional systems, whereas a rigorous framework for robust handling and conservative lower-bound schemes for the unavoidable energy or mode cutoffs when dealing with infinite-dimensional systems (such as quantum field theories) is detailed in Appendix F.
IV.1 Principle A: Quantum Hypothesis Testing
@@
From the fundamental perspective of information theory, the distinguishability between two quantum states is characterized by the theory of hypothesis testing. Theorem 3.2 has already revealed that the smoothed lower bound of RCC is directly related to the hypothesis testing relative entropy $D_{H}^{\eta}(\rho||\sigma_{R})$. This provides us with the first, and theoretically most fundamental, principle of information extraction. Its core task can be formulated as a binary quantum hypothesis testing problem: given a single-copy system, determine whether its state satisfies the null hypothesis $H_{0}:\sigma_{R}$ or the alternative hypothesis $H_{1}:\rho$. Any given measurement (described by a POVM $\{T,I-T\}$) will yield a lower bound on $D_{H}^{\eta}(\rho||\sigma_{R})$, which alleviates the stringent requirement of finding the optimal measurement, making this principle feasible in practice. At the implementation level, let us assume $N$ independent and repeated measurements are performed, from which the frequency of the type-II error (mistaking $\rho$ for $\sigma_{R}$) is observed, and its probability $\beta$ is thereby estimated. To ensure statistical rigor, we need to construct a confidence interval for $\beta$. Using standard statistical methods, such as the Hoeffding bound or the more precise Clopper-Pearson interval [Wilson1927, Hoeffding1963], one can calculate the upper confidence endpoint of $\beta$, denoted $\beta_{U}(\delta)$, at a confidence level of $1-\delta$. (Note: In the chosen binary test $\{T,I-T\}$, it is specified that observing $T$ leads to the decision $H_{1}(\rho)$, while observing ITI-T leads to the decision $H_{0}(\sigma_{R})$. The type-I error is constrained as $\alpha=\mathrm{Tr}[T\cdot\sigma_{R}]\leq\eta$. With a finite sample size, the upper confidence endpoint of $\alpha$, $\alpha_{U}(\delta)$, can be obtained using the Clopper–Pearson (or Wilson) interval, and it should be ensured that $\alpha_{U}(\delta)\leq\eta$; under this constraint, the lower bound of $D_{H}^{\eta}(\rho||\sigma_{R})$ is then assessed based on the upper confidence endpoint of $\beta$, $\beta_{U}(\delta)$.) We then have a statistical lower bound for the leading term of RCC:
@@
$D_{H}^{\eta}(\rho||\sigma_{R})\geq-\log\beta_{U}(\delta)$ (9)
This inequality explicitly relates the complexity of a state $\rho$ to its statistical distinguishability from the reference state $\sigma_{R}$ A highly complex state must have extremely high distinguishability, corresponding to an extremely low type-II error rate, which in turn implies a large lower bound on its complexity. The complete experimental protocol, including the rigorous control of statistical errors, is detailed in Appendix D.1. Of particular importance is that the advantage of the present theoretical framework lies in the fact that this connection between “distinguishability” and “complexity” does not merely remain at the level of physical intuition. Theorem 3.2 formalizes this connection as a rigorous mathematical inequality. It is this theorem, along with its complete derivation based on one-shot information theory in Appendix A.6, that builds the bridge from experimental data (the type-II error rate $\beta$) to the theoretical lower bound ($C_{\mathrm{opt}}$). This rigorous connection ensures that our proposed operational schemes are theoretically completely reliable and self-consistent, successfully transforming an abstract concept of complexity into a measurable physical reality.
IV.2 Principle B: Projective Witnesses
@@
$D_{\max}(\rho||\sigma_{R})\geq\log(p\cdot d_{R}/r)$ (10)
Proof. If $\rho\leq 2^{\lambda}\cdot\sigma_{R}$, then for a projector $\Pi$ of rank $r$, taking the trace yields $\mathrm{Tr}(\Pi\cdot\rho)\leq 2^{\lambda}\cdot\mathrm{Tr}(\Pi\cdot\sigma_{R})=2^{\lambda}\cdot r/d_{R}$. From $p=\mathrm{Tr}(\Pi\cdot\rho)$, one can solve for $\lambda\geq\log(p\cdot d_{R}/r)$. The lemma is thus proven. In experimental practice, we estimate the value of $p$ through $N$ measurements and calculate its lower confidence endpoint $p_{\mathrm{lower}}(\delta)$ at a confidence level of $1-\delta$. Substituting this statistical lower bound into Eq. (10) yields a lower bound for $D_{\max}$ with statistical confidence. The physical meaning of this result is exceptionally clear: if it is found experimentally that the quantum state $\rho$ occupies, with a high probability $p_{\mathrm{lower}}$, a subspace (defined by $\Pi$) of dimension $r$ much smaller than the dimension $d_{R}$ of its containing structured subspace, then this serves as a proof that the information in this state is highly “concentrated” or “ordered.” This high concentration of information implies that the state $\rho$ possesses a very large eigenvalue, and thus $H_{\mathrm{min}}(\rho)$ is small and $D_{\max}$ is large. Therefore, through a simple projective measurement, we have “witnessed” the non-triviality of the state, setting a rigorous lower bound on its complexity. The detailed implementation protocol for this principle, including the statistical estimation of occupation probabilities and confidence level corrections in scenarios with multiple witnesses, is formally articulated in Appendix D.2.
IV.3 Principle C: Dephasing Information Bound
The third principle extracts information through a “classicization” process. To obtain a reliable lower bound, we employ the completely dephasing channel $\Delta_{R}$ derived from the reference set $R$, which corresponds to performing a measurement in a basis of rank-1 projectors that simultaneously diagonalizes all projection operators $\{P_{i}\}$ and then discarding the off-diagonal (coherent) information. Physically, this channel corresponds to the strongest decoherence process that leaves the structure $R$ invariant.[HiaiPetz1991] According to the property of entropy increase in quantum information theory, the von Neumann entropy after the action of this channel is necessarily no less than before, i.e., $S(\Delta_{R}(\rho))\geq S(\rho)$. The dephased state $\Delta_{R}(\rho)$ is a classical probability distribution, and its Shannon entropy is denoted as $H=S(\Delta_{R}(\rho))$. Therefore, we have $H\geq S(\rho)$, which provides an operational and conservative lower bound for RCC:
@@ IV.3 Principle C: Dephasing Information Bound
Experimentally, this corresponds to performing measurements in a classical basis compatible with the reference set $R$, reconstructing the probability distribution, and estimating its Shannon entropy $H$. An upper confidence interval for $H$ can be given through the DKW inequality or bootstrap methods,[Paninski2003] thereby propagating the statistical uncertainty to the final lower bound of RCC. The detailed implementation protocol for this principle, particularly concerning how to rigorously estimate Shannon entropy from finite samples with bias correction, is presented with its complete statistical framework in Appendix D.3. Furthermore, the three principles proposed in this chapter can be integrated into an operational pipeline with statistical confidence, as detailed in Appendix B.
IV.4 An Important Clarification: The Role of Purity